Wolsey Industries Inc. expects to maintain the same inventories at the end of 20Y3 as at the beginning of the year. The total of all production costs for the year is therefore assumed to be equal to the cost of goods sold. With this in mind, the various department heads were asked to submit estimates of the costs for their departments during the year. A summary report of these estimates is as follows: Estimated Fixed Cost Variable Cost (per unit sold) 2 Production costs 3Direct materials 4 Direct labor 5 Factory overhead 6Selling expenses: $56.00 36.00 20.00 $194,000.00 Sales salaries and commissions 8.00
Answer #1

Wolsey Industries Inc

  1. Computation of expected contribution margin ratio:

Contribution margin ratio = contribution margin/sales

Contribution margin = sales – total variable costs

Estimated Sales = $3,424,000

Variable costs –

Direct materials$1,198,400

Direct labor$770,400

Variable OH$428,000

Salaries and commission $171,200

Misc. selling expense$21,400

Supplies expense$128,400

Misc. admin expense$21,400

Total variable costs = $ 2,739,200

Contribution margin = 3,424,000 – 2,739,200 = $684,800

Contribution margin per unit = $684,800/21,400 units = $32

Contribution margin ratio = 32/160 = 20%

Hence, expected contribution margin ratio = 20%

  1. Determination of break-even sales in units and dollars:

BEP in units = fixed cost/contribution margin per unit

Fixed cost –

Factory overhead$194,000

Sales salaries and commissions $110,000

Misc. selling expense$7,000

Admin. Salaries expense$124,600

Misc. admin expense$15,000

Total Fixed Costs$513,600

BEP in units = $513,600/$32 = 16,050 units

Break-even point sales in units = 16,050

= $513,600/20% = $2,568,000

BEP sales in dollars = $2,568,000

  1. Margin of safety in dollars and percentage

MOS in dollars = actual sales – BEP sales in dollars

= $3,424,000 - $2,568,000 = $856,000

MOS percentage = MOS/actual sales

= 856,000/3,424,000 = 25%

  1. Determination of operating leverage:

Operating leverage = contribution margin/net operating income

= $684,800/$171,200 = 4